# Debiasing Guidance for Discrete Diffusion with Sequential Monte Carlo

### v2

@@ 2.1 Continuous Time Markov chains
where $\delta_{a,b}$ is $1$ if $a=b$ and 00 otherwise.
@@ 3 Guidance
We defer the learning of guided rate matrix to Section 6.1. One might expect the guided rate matrix $R_{t}^{\alpha}$ to represent the time-reversal corrupting process starting at the tempered distribution $p_{0}(x_{0})p_{0}(\zeta|x_{0})^{\alpha}/{\mathcal{Z}}_{\alpha}$. However, as shown in Proposition 3.2, this is not the case. {mdframed}[style=highlightedBox]
@@
Let $M_{t}[\cdot]$ denote the evolution of probability mass at time $t$ under the corrupting process $p^{\text{corrupt}}_{t|0}$ of $X_{t}$, the unconditional diffusion model. Define $p_{t}^{\alpha,\text{true}}$ as:
Then,
The true tempered rate matrix $R_{t}^{\alpha,\text{true}}$ given by
satisfies the reverse-time Kolmogorov Forward Equation:
$\partial_{t}p_{t}^{\alpha,\text{true}}=-\left[R_{t}^{\alpha,\text{true}}\right]^{\intercal}p_{t}^{\alpha,\text{true}}.$
From Proposition 3.2, it follows that $R_{t}^{\alpha}=R_{t}^{\alpha,\text{true}}$ holds only for $\alpha=1$. Furthermore, for $\alpha=1$, guidance requires that the base distribution $p_{1}$ is independent of the conditioning variable $\zeta$ to correctly sample from the conditioned distribution $p_{0}(X_{0}|\zeta)$, as shown in Corollary 3.3.
@@ 4.1 Importance Sampling the Tempered Distribution
Let proposal $\{Y_{t}\}_{t\in[0,1]}$ be a reverse-time CTMC with rate matrix $Q_{t}$ and initial distribution $p_{1}(\cdot)$. Define $\{W_{t}\}_{t\in[0,1]}$ by
$W_{t}=\frac{\mathrm{d}\mathbb{P}_{t}^{\mathrm{base}}}{\mathrm{d}\mathbb{Q}_{t}}\left(\frac{\mathrm{d}\mathbb{P}^{\alpha=1}_{t}}{\mathrm{d}\mathbb{P}_{t}^{\mathrm{base}}}\right)^{\alpha},$
where we have path space measures ${\mathbb{P}}_{t}^{\text{base}}=\mathrm{Law}{\{X_{\tau}\}_{\tau\in[t,1]}}$, ${\mathbb{P}}^{\alpha=1}_{t}=\mathrm{Law}{\{X_{\tau}\mid\zeta\}_{\tau\in[t,1]}}$ and ${\mathbb{Q}}_{t}=\mathrm{Law}{\{Y_{\tau}\}_{\tau\in[t,1]}}$.
Then the Radon-Nikodym derivatives can be written as,
$\ln\frac{\mathrm{d}\mathbb{P}_{t}}{\mathrm{d}\mathbb{Q}_{t}}^{\!\mathrm{base}}=$ $\displaystyle\sum_{\begin{subarray}{c}\tau\geq t\\ Y_{\tau^{+}}\neq Y_{\tau}\end{subarray}}\!\!\ln R_{t}(Y_{\tau^{+}},Y_{\tau})$
$+\int_{1}^{t}Q_{\tau}(Y_{t})-R_{\tau}(Y_{t})\,\mathrm{d}\tau$
and
$\ln\!\frac{\mathrm{d}\mathbb{P}_{t}^{\alpha=1}}{\mathrm{d}\mathbb{P}_{t}^{\mathrm{base}}}\!=$ $\!\!\!\!\!\displaystyle\sum_{\begin{subarray}{c}\tau\geq t\\ Y_{\tau^{+}}\neq Y_{\tau}\end{subarray}}\!\!\!\!\ln R_{t}^{\alpha=1}\!(Y_{\tau^{+}},Y_{\tau}|\zeta)$
$+\!\!\int_{1}^{t}\!\!\!R_{\tau}(Y_{\tau})-R^{\alpha=1}_{\tau}(Y_{\tau}|\zeta)\mathrm{d}\tau,$
such that for any $p_{t}^{\alpha}$-integrable function $h$ and time $t\in[0,1]$, we have
$\mathbb{E}_{x\sim p_{t}^{\alpha}}[h(x)]=\frac{\mathbb{E}[W_{t}\cdot h(Y_{t})]}{\mathbb{E}[W_{t}]},$
where the expectation is taken over $\text{Law}(W_{t},Y_{t})$
@@
For practical implementation, we approximate the CTMC by discretising time into $T$ steps: $01=t_{1}>t_{2}>\dots>t_{T}=$ 0. For times $t<s$, we denote the transitions derived from rate matrices $Q_{t}$, $R_{t}$, and $R^{\alpha=1}_{t}$ as $q_{t|s}(x_{t}|x_{s})$, $p_{t|s}(x_{t}|x_{s})$, and $p_{t|s}(x_{t}|x_{s},\zeta)$ respectively. To this end, we consider a discretisation of Proposition 4.1 in Appendix E. We also present Algorithm 1 in Section 4.2, a pseudo-code of the discretised version of the proposed sampling method.
@@ 4.2 Resampling and Sequential Monte Carlo
Sequential Monte Carlo (SMC) addresses this challenge by performing resampling steps at intermediate times $10<t<$ 1, obtaining approximate samples of $p_{t}^{\alpha}$ while resetting the weights to maintain the importance sampling equality. Effectively, this eliminates sample from low likelihood regions and duplicates those in high likelihood regions.
@@ 6.4.2 Ablating Guidance and SMC temperature
Perplexity $65.048$±$\pm 30.951$ $67.062$±$\pm 42.959$ $73.562$±$\pm 26.980$ $80.583$±$\pm 45.478$ $63.915$±$\pm 31.375$ $53.697$±$\pm 26.543$
BERTScore $\uparrow$ $0.583$±$\pm 0.038$ $0.584$±$\pm 0.032$ $0.587$±$\pm 0.029$ $0.576$±$\pm 0.038$ $0.583$±$\pm 0.037$ $0.595$±$\pm 0.038$
GLEU $\uparrow$ $0.701$±$\pm 0.021$ $0.703$±$\pm 0.022$ $0.708$±$\pm 0.019$ $0.699$±$\pm 0.022$ $0.701$±$\pm 0.021$ $0.706$±$\pm 0.021$
Accuracy $\uparrow$ $0.992$±$\pm 0.015$ $0.996$±$\pm 0.014$ $0.998$±$\pm 0.013$ $0.998$±$\pm 0.015$ $0.991$±$\pm 0.018$ $1.000$±$\pm 0.000$
Table 4: Performance of guidance methods on text infilling: Sequential Monte Carlo (varying twist temperature $\alpha$) plus a row of guidance temperature (all set to 1.0), Guidance, and Reconstruction guidance. Values show mean $\pm$ standard deviation. $\uparrow$/$\downarrow$ indicate better performance for higher/lower values. Light blue cell marks the best result. Light green cell marks the best result between Guidance and SMC.
@@ 6.4.3 Scaling up Discretisation Step
Table 4, Table 2, and Table 3 summarise our empirical results. SMC with a high SMC temperature significantly outperforms the base guidance method across all tasks. Notably, for toxicity and sentiment controlled generation, our SMC-based method significantly outperforms a simple prompting strategy in generating strong conditions. However, for infilling and despite their theoretical guarantees, guidance and SMC approaches underperform compared to the reconstruction method, even though it is known to be biased (Wu et al., 2024). This indicates an inadequate learning of the guided rate matrix. While SMC can strengthen the base guidance method, it cannot fully compensate for an imperfectly learned guided rate matrix.
@@ B.2 The Transition Rate Matrix
where $\delta_{a,b}$ is 1 if $a=b$ and 00 if otherwise.
@@ Appendix C Proof for Section 3
satisfies the reverse-time Kolmogorov Forward Equation:
$\partial_{t}p_{t}^{\alpha,\text{true}}=-\left[R_{t}^{\alpha,\text{true}}\right]^{\intercal}p_{t}^{\alpha,\text{true}}.$
@@
We denote $Z_{t,\alpha}$ the normalising constant of $p_{t}^{\alpha}$ such that:
$p_{t}^{\alpha,\text{true}}(x_{t})=\frac{1}{Z_{t,\alpha}}p_{t}(x_{t})\mathbb{E}[p(\zeta|X_{0})^{\alpha}|X_{t}=x_{t}]$ (31)
An other intermediate result that we need is to show that the normalising constant $Z_{t,\alpha}$ is independent of the time $t$
$Z_{t,\alpha}$ $=\sum_{y}p_{t}^{\alpha}(y)$ (32)
$=\sum_{y}p_{t}(y)\mathbb{E}[p(\zeta|X_{0})^{\alpha}|X_{t}=y]$ (33)
$=\sum_{y}\sum_{x_{0}}p_{t}(y)p(\zeta|x_{0})^{\alpha}p_{0|t}(x_{0}|y)$ (34)
$=\sum_{y}\sum_{x_{0}}p(\zeta|x_{0})^{\alpha}p_{t|0}(y|x_{o})p_{0}(x_{0})$ (35)
$=\sum_{x_{0}}p(\zeta|x_{0})^{\alpha}p_{0}(x_{0})\left(\sum_{y}p_{t|0}(y|x_{0})\right)$ (36)
$=\sum_{x_{0}}p(\zeta|x_{0})^{\alpha}p_{0}(x_{0})$ (37)
$=\mathbb{E}[p(\zeta|X_{0})^{\alpha}]$ (38)
We see that $Z_{t,\alpha}$ equals $\mathbb{E}[p(\zeta|X_{0})^{\alpha}]$ which is independent of $t$ and as a consequence $\partial_{t}Z_{t,\alpha}=0$. We also denote the normalising constant as $Z_{\alpha}:=Z_{t,\alpha}$. We now prove the main result of the proposition with the following derivation:
We first introduce the following notation to make derivation lighter:
$w_{t,\alpha}(x_{t})=\mathbb{E}[p(\zeta|X_{0})^{\alpha}|X_{t}=x_{t}]$ (39)
Noticing $w$ satisfies a reverse-time Kolmogorov Backward Equation, we have:
$\partial_{t}p_{t}^{\alpha}(x_{t})$ $=\partial_{t}[\frac{1}{Z_{\alpha}}p_{t}(x_{t})w_{t,\alpha}(x_{t})]$ (40)
$=\frac{1}{Z_{\alpha}}\left[p_{t}(x_{t})\partial_{t}w_{t,\alpha}(x_{t})+w_{t,\alpha}(x_{t})\partial_{t}p_{t}(x_{t})\right]$ (41)
$=\frac{1}{Z_{\alpha}}\left[\sum_{y}w_{t,\alpha}(y)p_{t}(x_{t})R_{t}(x_{t},y)-\sum_{y}w_{t,\alpha}(x_{t})p_{t}(y)R_{t}(y,x_{t})\right]$ (42)
$=\frac{1}{Z_{\alpha}}\left[\sum_{y\neq x_{t}}w_{t,\alpha}(y)p_{t}(x_{t})R_{t}(x_{t},y)-\sum_{y\neq x_{t}}w_{t,\alpha}(x_{t})p_{t}(y)R_{t}(y,x_{t})\right]$ (43)
$=\frac{1}{Z_{\alpha}}\left[\sum_{y\neq x_{t}}p_{t}(x_{t})w_{t,\alpha}(x_{t})R_{t}(x_{t},y)\frac{w_{t,\alpha}(y)}{w_{t,\alpha}(x_{t})}-\sum_{y\neq x_{t}}p_{t}(y)w_{t,\alpha}(y)R_{t}(y,x_{t})\frac{w_{t,\alpha}(x_{t})}{w_{t,\alpha}(y)}\right]$ (44)
$=\sum_{y\neq x_{t}}p_{t}^{\alpha,\text{true}}(x_{t})R^{\alpha}_{t}(x_{t},y)-\sum_{y\neq x_{t}}p_{t}^{\alpha,\text{true}}(y)R^{\alpha}_{t}(y,x_{t})$ (45)
and the last equation is the reverse time Forward Kolmogorov Equation which concludes the proof.
@@ D.1 Itô Process Counterpart
$\ln\!\frac{\mathrm{d}\mathbb{P}_{t}^{\alpha=1}}{\mathrm{d}\mathbb{P}_{t}^{\mathrm{base}}}\!=\int_{t}^{1}\frac{1}{g_{\tau}^{2}}(f_{\tau}^{\cdot|\zeta}(Y_{\tau})-f_{\tau}(Y_{\tau}))^{\top}\overleftarrow{{\mathrm{d}}}Y_{\tau}+\int_{t}^{1}\frac{1}{2g_{\tau}^{2}}(||f^{\cdot|\zeta}_{\tau}(Y_{\tau})||^{2}-||f_{\tau}(Y_{\tau})||^{2}){\mathrm{d}}\tau$
@@
[…]
@@
[…]
@@
$=R^{\alpha=1}_{t}(x_{t})-R_{t}(x_{t})$ Definition of $\text{Definition of }R_{t}^{\alpha=1}$ (75)
@@ Appendix E Statement and Derivation of Discretised Weights
$=\frac{1}{{\mathcal{Z}}_{1}^{\alpha}}\sum_{Y_{t_{k}}}h(Y_{t_{k}})p_{t}(\zeta|Y_{t_{k}})^{\alpha}p_{t}(Y_{k})$ Both $\text{Both }Y_{t}\text{ and }X_{t}\text{ starts at }p_{1}$

### v3

@@ 2.1 Continuous Time Markov chains
where $\delta_{a,b}$ is $1$ if $a=b$ and 0 otherwise.
@@ 3 Guidance
We defer the learning of guided rate matrix to Section 6.1. One might expect the guided rate matrix $R_{t}^{\alpha}$ to represent the time-reversal corrupting process starting at the tempered distribution $p_{0}(x_{0})p_{0}(\zeta|x_{0})^{\alpha}/{\mathcal{Z}}_{\alpha}$. However, as shown in Proposition 3.2, this is not the case.
@@
Let $M_{t}[\cdot]$ denote the evolution of probability mass at time $t$ under the corrupting process $p^{\text{corrupt}}_{t|0}$ of $X_{t}$, the unconditional diffusion model. Define $p_{t}^{\alpha,\text{true}}$ as: $p_{t}^{\alpha,\text{true}}=M_{t}\big{[}p_{0}(\cdot)\,p_{0}(\zeta|\cdot)^{\alpha}/{\mathcal{Z}}_{\alpha}\big{]}.$ Then, $p_{t}^{\alpha,\text{true}}(x_{t})\propto p_{t}(x_{t})\mathbb{E}\big{[}p_{0}(\zeta|X_{0})^{\alpha}|X_{t}=x_{t}\big{]}.$ The true tempered rate matrix $R_{t}^{\alpha,\text{true}}$ given by $\forall x\neq y,\;R_{t}^{\alpha,\text{true}}(x,y|\zeta)$ $=R_{t}(x,y)\tfrac{\mathbb{E}[p_{0}(\zeta|X_{0})^{\alpha}|X_{t}=y]}{\mathbb{E}[p_{0}(\zeta|X_{0})^{\alpha}|X_{t}=x]},$ $\forall x,\;R_{t}^{\alpha,\text{true}}(x,x|\zeta)$ $=-\sum_{y\neq x}R_{t}^{\alpha,\text{true}}(x,y|\zeta),$ is rate matrix of time reversal of the corruption process starting from $p_{0}(\cdot)p_{0}(\zeta|\cdot)^{\alpha}/{\mathcal{Z}}_{\alpha}$ From Proposition 3.2, it follows that $R_{t}^{\alpha}=R_{t}^{\alpha,\text{true}}$ holds only for $\alpha=1$. Furthermore, for $\alpha=1$, guidance requires that the base distribution $p_{1}$ is independent of the conditioning variable $\zeta$ to correctly sample from the conditioned distribution $p_{0}(X_{0}|\zeta)$, as shown in Corollary 3.3.
@@ 4.1 Importance Sampling the Tempered Distribution
Let proposal $\{Y_{t}\}_{t\in[0,1]}$ be a reverse-time CTMC with rate matrix $Q_{t}$ and initial distribution $p_{1}(\cdot)$. Define $\{W_{t}\}_{t\in[0,1]}$ by $W_{t}=\frac{\mathrm{d}\mathbb{P}_{t}^{\mathrm{base}}}{\mathrm{d}\mathbb{Q}_{t}}\left(\frac{\mathrm{d}\mathbb{P}^{\alpha=1}_{t}}{\mathrm{d}\mathbb{P}_{t}^{\mathrm{base}}}\right)^{\alpha},$ where we have path space measures ${\mathbb{P}}_{t}^{\text{base}}=\mathrm{Law}{\{X_{\tau}\}_{\tau\in[t,1]}}$, ${\mathbb{P}}^{\alpha=1}_{t}=\mathrm{Law}{\{X_{\tau}\mid\zeta\}_{\tau\in[t,1]}}$ and ${\mathbb{Q}}_{t}=\mathrm{Law}{\{Y_{\tau}\}_{\tau\in[t,1]}}$. Then the Radon-Nikodym derivatives can be written as, $\ln\frac{\mathrm{d}\mathbb{P}_{t}}{\mathrm{d}\mathbb{Q}_{t}}^{\!\mathrm{base}}=$ $\displaystyle\sum_{\begin{subarray}{c}\tau\geq t\\ Y_{\tau^{+}}\neq Y_{\tau}\end{subarray}}\!\!\ln R_{t}(Y_{\tau^{+}},Y_{\tau})$ $+\int_{1}^{t}Q_{\tau}(Y_{t})-R_{\tau}(Y_{t})\,\mathrm{d}\tau$ and $\ln\!\frac{\mathrm{d}\mathbb{P}_{t}^{\alpha=1}}{\mathrm{d}\mathbb{P}_{t}^{\mathrm{base}}}\!=$ $\!\!\!\!\!\displaystyle\sum_{\begin{subarray}{c}\tau\geq t\\ Y_{\tau^{+}}\neq Y_{\tau}\end{subarray}}\!\!\!\!\ln R_{t}^{\alpha=1}\!(Y_{\tau^{+}},Y_{\tau}|\zeta)$ $+\!\!\int_{1}^{t}\!\!\!R_{\tau}(Y_{\tau})-R^{\alpha=1}_{\tau}(Y_{\tau}|\zeta)\mathrm{d}\tau,$ such that for any $p_{t}^{\alpha}$-integrable function $h$ and time $t\in[0,1]$, we have $\mathbb{E}_{x\sim p_{t}^{\alpha}}[h(x)]=\frac{\mathbb{E}[W_{t}\cdot h(Y_{t})]}{\mathbb{E}[W_{t}]},$ where the expectation is taken over $\text{Law}(W_{t},Y_{t})$
@@
For practical implementation, we approximate the CTMC by discretising time into $T$ steps: $1=t_{1}>t_{2}>\dots>t_{T}=0$. For times $t<s$, we denote the transitions derived from rate matrices $Q_{t}$, $R_{t}$, and $R^{\alpha=1}_{t}$ as $q_{t|s}(x_{t}|x_{s})$, $p_{t|s}(x_{t}|x_{s})$, and $p_{t|s}(x_{t}|x_{s},\zeta)$ respectively. To this end, we consider a discretisation of Proposition 4.1 in Appendix E. We also present Algorithm 1 in Section 4.2, a pseudo-code of the discretised version of the proposed sampling method.
@@ 4.2 Resampling and Sequential Monte Carlo
Sequential Monte Carlo (SMC) addresses this challenge by performing resampling steps at intermediate times $0<t<1$, obtaining approximate samples of $p_{t}^{\alpha}$ while resetting the weights to maintain the importance sampling equality. Effectively, this eliminates sample from low likelihood regions and duplicates those in high likelihood regions.
@@ 6.4.2 Ablating Guidance and SMC temperature
[…]
Table 4: Performance of guidance methods on text infilling: Sequential Monte Carlo (varying twist temperature $\alpha$) plus a row of guidance temperature (all set to 1.0), Guidance, and Reconstruction guidance. Values show mean $\pm$ standard deviation. $\uparrow$$\downarrow$ indicate better performance for higher/lower values. Light blue cell marks the best result. Light green cell marks the best result between Guidance and SMC.
@@ 6.4.3 Scaling up Discretisation Step
Table 4, Table 2, and Table 3 summarise our empirical results. SMC with a high SMC temperature significantly outperforms the base guidance method across all tasks. Notably, for toxicity and sentiment-controlled generation, our SMC-based method significantly outperforms a simple prompting strategy in generating strong conditions. However, for infilling and despite their theoretical guarantees, guidance and SMC approaches underperform compared to the reconstruction method, even though it is known to be biased (Wu et al., 2024). This indicates an inadequate learning of the guided rate matrix. While SMC can strengthen the base guidance method, it cannot fully compensate for an imperfectly learned guided rate matrix.
@@ 7 Conclusion and Future Work
Acknowledgement
The authors would like to thank Junhua Chen for extensive dicussions on the implementation of Sequential Monte Carlo Methods.
@@ B.2 The Transition Rate Matrix
where $\delta_{a,b}$ is 1 if $a=b$ and 0 if otherwise.
@@ Appendix C Proof for Section 3
is time-reversal of the corruption process that stats from $p_{0}(\cdot)p_{0}(\zeta|\cdot)^{\alpha}/{\mathcal{Z}}_{\alpha}$
@@
Then it follow from standard results of time-reversal of the CTMC such as that in (Lou et al., 2024), the rate matrix of the time-reversal can be written as:
$\forall x\neq y.\quad R_{t}^{\alpha,\text{true}}(x,y|\zeta)=R_{t}(x,y)\frac{p_{t}^{\alpha,\text{true}}(y)}{p_{t}^{\alpha,\text{true}}(x)}=R_{t}^{\alpha,\text{true}}(x,y|\zeta)$ $=R_{t}(x,y)\tfrac{\mathbb{E}[p_{0}(\zeta|X_{0})^{\alpha}|X_{t}=y]}{\mathbb{E}[p_{0}(\zeta|X_{0})^{\alpha}|X_{t}=x]}$ (31)
@@ D.1 Itô Process Counterpart
[…]
@@
Define $g_{t}(y)=\int_{{\mathbb{R}}}e^{a}f_{t}(y,a)da$. Then, via dominated convergence theorem (i.e. $\partial_{t}g_{t}(y)=\int_{{\mathbb{R}}}e^{a}\partial_{t}f_{t}(y,a)da$) and substituting $\partial_{t}f_{t}(y,a)$ from Equation 40, it follows that
@@
And finally, by integration by parts, and noting that $\int_{\mathbb{R}}e^{a}f_{t}(y,a)da=Q_{t}(y)g_{t}(y)$, we have
@@
$=R^{\alpha=1}_{t}(x_{t})-R_{t}(x_{t})$ $\text{Definition of }R_{t}^{\alpha=1}$ (60)
@@ Appendix E Statement and Derivation of Discretised Weights
$=\frac{1}{{\mathcal{Z}}_{1}^{\alpha}}\sum_{Y_{t_{k}}}h(Y_{t_{k}})p_{t}(\zeta|Y_{t_{k}})^{\alpha}p_{t}(Y_{k})$ $\text{Both }Y_{t}\text{ and }X_{t}\text{ starts at }p_{1}$